### v3

# Information About Other Agents in Mechanism Design

Abstract
We study mechanism design settings where the planner has an interest in agents receiving noisy signals about the types of other agents. When the planner is interested only in partial implementation — designing a mechanism for which there exists an equilibrium that yields a desired outcome — any social choice rule that is incentive-compatible after agents receive additional information about other agents is incentive-compatible without this information. However, additional information about other agents can eliminate undesired equilibria, making it helpful to a planner interested in full implementation — designing a mechanism for which every equilibrium outcome is desirable. We provide a sufficient condition under which a social choice function that is not fully implementable when agents have no information about types of other agents can become fully implementable if agents have additional information.
1 Introduction
The goal of a social planner in any mechanism design problem is to design a game under which a desirable outcome emerges in equilibrium. Agents have private information about their own preferences and preferences of other agents. Naturally, the set of equilibrium outcomes that can emerge from a mechanism is sensitive to the details of the information available to agents. A considerable body of mechanism design literature focuses on the effects of changing information structures on achieving desirable outcomes. Most of this work focuses on how much agents know about variables that directly affect their own payoffs — for example, the information a bidder knows about her own valuation for an object (Bergemann et al. (2022)) — rather than information about other agents’ preferences.
In this paper, we examine how the planner can have an interest in agents possessing additional information about the preferences of other agents, revisiting the Bayesian implementation problem as studied by Jackson (1991) (see also Postlewaite and Schmeidler (1986) and Palfrey and Srivastava (1989)). In this setting, agents fully understand their own preferences over a set of outcomes but do not know the preferences of other agents. We consider how the implementability of a social choice function (SCF) changes if agents can receive a full-support signal regarding the preferences of other agents. In particular, our main result gives a sufficient condition under which a planner interested in full implementation — designing a mechanism such that every equilibrium yields a desirable outcome — will strictly prefer an agent to have additional information about agents.
In section 3, we show that any SCF which satisfies incentive compatibility when agents receive informative signals about the types of other agents also satisfies incentive compatibility when agents do not have this additional information. This implies that a planner concerned only with partial implementation — ensuring that the desirable outcome constitutes an equilibrium of the mechanism — will never strictly prefer an agent to have additional information about other agents. The argument is straightforward: When an agent has access to an informative signal about the preferences of other agents, the planner must ensure that the agent is incentivized to report truthfully in each signal realization. A weighted average of these constraints across signal realizations yields the incentive constraint needed for the agent if she did not have access to additional information, making this latter constraint easier to satisfy.
The remainder of the paper examines how additional information about the types of other agents can help an agent signal untruthful equilibria to the planner. In section 4, we present the main example of the paper: an SCF that is not fully implementable when agents have no information about the preferences of other agents but is fully implementable when agents have more information. The condition that rules out undesired equilibria in Bayesian implementation settings is Bayesian monotonicity, which guarantees that agents have sufficient incentives to signal to the planner that an undesirable equilibrium strategy is being played. We focus heavily on how Bayesian monotonicity varies with the information agents have about other agents. Section 5 presents a sufficient condition for an SCF that is not fully implementable when agents have no information about the types of other agents to be fully implementable when agents have more information.
An obvious concern with giving an agent additional information about types of other agents is that it can affect their own incentive to report truthfully to the mechanism. Indeed, the conclusion of section 3 is that incentive compatibility becomes (weakly) harder to satisfy if agents receive additional information about other agents. To avoid eliminating truthful reporting as an equilibrium, we only modify the information available to agents for which truth-telling is a dominant strategy in the direct mechanism of a given SCF. Our results do not speak to SCFs for which no agent has a dominant strategy in the direct mechanism.
To establish our main result, we note that an incentive-compatible SCF is not fully implementable if and only if there is an untruthful equilibrium of the direct mechanism such that no agent has an incentive to report the deception to the planner. The task is then to construct a signal that gives some agent $i$ the ability to eliminate these equilibria. In general, if truth-telling is a dominant strategy for player $i$, an untruthful equilibrium $\alpha$ of the direct mechanism will involve some subset of other agents playing deceptions. The signal we construct partitions the set of other agents’ types into two groups. Each group in the partition must contain a type profile of other agents such that if the agent $i$ knew this type profile realized, she would prefer all other agents to report truthfully rather than deceive according to $\alpha$, regardless of her own type. If agent $i$ is confident enough that the realized type profile is in one group or the other, she can propose an alternate SCF that swaps the outcomes implemented under the truthful strategy versus the deception. By construction, this alternate SCF can only give agent $i$ a weakly lower payoff relative to the original if all other agents report truthfully, so agent $i$ would only propose it if $\alpha$ is being played. Whenever the signal is sufficiently accurate, agent $i$ can eliminate $\alpha$ as an equilibrium. We conclude by showing that this procedure can iteratively eliminate all untruthful equilibria of the direct mechanism.
Though a planner may want agents to have access to informative signals about the types of other agents compared to no information, our analysis also demonstrates that a planner may prefer an information structure that lies strictly between no information and complete information. This conclusion is consistent with the rapidly growing literature on information design spurred by Rayo and Segal (2010) and Kamenica and Gentzkow (2011).
Related Literature
This paper directly contributes to the literature on full Bayesian implementation — which began with Postlewaite and Schmeidler (1986), Palfrey and Srivastava (1989), and Jackson (1991) — by shedding light on the relationship between Bayesian monotonicity and the information available to agents. Since agents are receiving informative signals about the state of the world, we also add to the extensive literature on the effects of changing information structures in mechanism design settings pioneered by Bergemann and Välimäki (2002). It relates especially to the literature on information disclosure in mechanism design. In auction settings, Ganuza (2004) studies the optimal way for a seller to disclose a costly public signal about a good’s characteristic in a second-price auction, which is a component of bidders’ private valuations, while Bergemann and Pesendorfer (2007) investigate the joint design problem in which the seller controls the information bidders have about their own valuations and the auction rules. Other papers that have investigated the effects of information disclosure in auction settings include Eső and Szentes (2007), and Bergemann et al. (2022). Similar questions have also been investigated in screening settings by Li and Shi (2017) and Roesler and Szentes (2017). Since much of the existing literature studies cases when agents’ types are independently drawn, existing results focus on how much agents know about their own preferences. This paper focuses on a mechanism designer’s interest in agents’ information about types of other agents. Additionally, unlike the auction papers that portray the auctioneer as having the ability to provide information to agents, this paper remains agnostic on where the information about other agents comes from. The emphasis is on how changing information structures can affect the planner’s implementation goals. Finally, while our paper asks whether a planner can be interested in agents’ information about other agents, one can also ask about an agent’s incentives to acquire such information, as done in Gleyze and Pernoud (2023), who show that most mechanisms incentivize agents to acquire information about other agents in environments where agents have uncertainty about their own types as well as types of other agents.
2 Model
[…]
The planner wishes to implement a social choice function (SCF) $F:\Theta\to A.$ The planner commits to a mechanism $\Gamma=\langle M,h\rangle$, such that $M=\prod_{i=1}^{n}M_{i}$ consists of a message space $M_{i}$ for each agent $i$ and $h:M\to A$ is the outcome function. A strategy profile $s=(s_{1},\dots,s_{n})$ specifies for each agent $i$ a function $s_{i}:\Theta_{i}\to\Delta(M_{i})$. At any state $\theta$, let $s(\theta)=(s_{1}(\theta_{1}),\dots,s(\theta_{n}))$ be the message profile sent by all agents, and $s_{-i}(\theta_{-i})=(s_{j}(\theta_{j}))_{j\neq i}$ denote the messages sent by agents other than $i.$ Let $B^{\Gamma}$ be the set of Bayesian Nash equilibria for a mechanism $\Gamma.$ If there is a $s\in B^{\Gamma}$ such that $h\circ s=F$ then $\Gamma$ partially implements $F$, and $F$ is considered partially implementable. If $\Gamma$ partially implements $F$, and $B^{\Gamma}$ contains only one equilibrium outcome, then $\Gamma$ fully implements $F$ and $F$ is considered fully implementable.
Classical mechanism design results pin down two key conditions for the implementability of an SCF $F$ in Bayesian settings.
Definition 1.
An SCF $F$ is (Bayesian) incentive-compatible if and only if for all agents $i$ and all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i},$
$\mathbb{E}(u_{i}(F(\theta_{i},\theta_{-i}),\theta_{i})\mid\theta_{i})\geq\mathbb{E}(u_{i}(F(\theta_{i}^{\prime},\theta_{-i}),\theta_{i})\mid\theta_{i}),$
where the expectation is taken over the belief of type $\theta_{i}$ about $\Theta_{-i},$ which is formed from the common prior $\mu$.
From the revelation principle we know that an SCF $F$ is partially implementable iff $F$ is incentive-compatible. For an incentive-compatible $F$ to be fully implementable, it must also satisfy Bayesian monotonicity, as studied by Postlewaite and Schmeidler (1986); Palfrey and Srivastava (1989); Jackson (1991).
Definition 2.
A deception for agent $i$ is a function $\alpha_{i}:\Theta_{i}\to\Theta_{i}.$ A deception $\alpha=(\alpha_{1},\dots,\alpha_{n})$ consists of a deception $\alpha_{i}$ for each agent $i\in N.$
Definition 3.
Let $\alpha$ be a deception and
$\alpha(\theta)=(\alpha_{1}(\theta_{1}),\dots,\alpha_{n}(\theta_{n}))$
and
$\alpha_{-i}(\theta_{-i})=(\alpha_{1}(\theta_{1}),\dots,\alpha_{i-1}(\theta_{i-1}),\alpha_{i+1}(\theta_{i+1}),\dots,\alpha_{n}(\theta_{n})).$
An SCF $F$ satisfies Bayesian monotonicity if for all deceptions $\alpha$ such that $F\circ\alpha\neq F$ there is an agent $i$ and a function $y:\Theta_{-i}\to A$ such that
$\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}),\theta_{i})\mid\theta_{i})\geq\mathbb{E}(u_{i}(y(\theta_{-i}),\theta_{i})\mid\theta_{i})$
for all $\theta_{i}\in\Theta_{i}$ and
$\mathbb{E}(u_{i}(F(\alpha(\theta_{-i},\theta_{i}^{\prime})),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})<\mathbb{E}(u_{i}(y(\alpha_{-i}(\theta_{-i})),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})$
for some $\theta_{i}^{\prime}\in\Theta_{i}.$
It was shown in Jackson (1991) that incentive compatibility and Bayesian monotonicity are necessary and sufficient for an SCF $F$ to be fully implementable in economic environments with at least 3 agents. [^2] An environment is economic if given any state $\theta\in\Theta$ and SCF $F$, at least two agents prefer some outcome $a\in A$ to $F(\theta).$ [^3] Jackson (1991) also identifies a closure condition for collections of SCFs, social choice sets, but the condition is vacuous for the case of a single-valued SCF. While incentive compatibility guarantees that truth-telling is an equilibrium of the direct mechanism, Bayesian monotonicity is the condition that allows the planner to design a mechanism that can eliminate undesired equilibria. The intuition behind the condition, which we will reference throughout this paper, is as follows. Suppose agents are initially at an equilibrium that uses a deception $\alpha$ such that $F\circ\alpha\neq F.$ By allowing any individual agent $i$ to specify an alternate SCF (the function $y$ can be extended to $\Theta$ by implementing the same outcome regardless of agent $i$’s type) that is implemented only if it appears to harm agent $i$, an agent can choose to use this alternate SCF to signal to the planner that the agents are lying. Upon seeing the alternate SCF, the planner understands that a deception is being played because it would harm agent $i$ in a truth-telling equilibrium, regardless of her own type. Bayesian monotonicity gives at least one type of agent $i$ an incentive to specify an alternate SCF because she would prefer the alternate SCF to $F$ under the deception $\alpha$, ruling out the deception as an equilibrium. Bayesian monotonicity guarantees that such a procedure can be done for every undesirable deception.
The beliefs agents have regarding the types of other agents is an important component of both incentive compatibility and Bayesian monotonicity. We study how these conditions change when agents have additional information about the types of other agents.
Suppose each agent $i$ can access a signal
$\sigma_{i}:\Theta_{-i}\to\Delta(\zeta_{i}),$
where $\zeta_{i}$ is a finite set of signal realizations and $\sigma_{i}(\theta_{-i})$ is the signal observed by agent $i$ when the type profile of other agents is $\theta_{-i}.$ Assume that $\sigma_{i}$ has full support over $\zeta_{i}$ for every agent $i.$ After observing the realization of $\sigma_{i}$ and her type $\theta_{i}$, agent $i$ updates her beliefs regarding $\Theta_{-i}$ (and $\Theta$) according to Bayes’ rule. In particular, suppose $\zeta_{i}=\{A_{i1},\dots,A_{ik}\}$ and agent $i$ has type $\theta_{i}.$ Upon observing $\sigma_{i}=A_{ij}$, agent $i$ places probability
$\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=A_{ij})=\frac{\mu(\theta_{-i}\mid\theta_{i})\sigma_{i}(\theta_{-i})(A_{ij})}{\sum_{\theta_{-i}^{\prime}}{\mu(\theta_{-i}^{\prime}\mid\theta_{i})\sigma_{i}(\theta_{-i}^{\prime})(A_{ij})}}$
on type $\theta_{-i}.$
The timing of the game is as follows.
Nature chooses $\theta$ from $\mu$, planner announces $\Gamma$, Structure of signals$\{\sigma_{i}\}_{i\in I}$ announcedAgents observe realizations of $\sigma_{i}$gents choose strategies $s_{i}$gents learn $\theta_{i}$,send messagesaccording to $s_{i}$utcome realized
Figure 1: Mechanism timing.
When the signals $\{\sigma_{i}\}_{i\in I}$ provide no information, the model is equivalent to the classical Bayesian implementation problem studied by Jackson (1991). The main question we address is whether and when an SCF $F$ that is not partially (resp. fully) implementable when $\sigma_{i}$ is noninformative for all $i$ can become partially (resp. fully) if one or more of the $\sigma_{i}$ is nontrivial. We are agnostic on where the signals $\{\sigma_{i}\}_{i\in I}$ come from. The nature of the question is one of comparative statics on the implementability of an SCF $F$ with respect to an increase in information one more agents $i$ has regarding $\Theta_{-i}$. [^4] Since agents initially do not have any information about $\Theta_{-i}$ aside from the common prior $\mu$, any nontrivial signal $\sigma_{i}$ represents an increase in agent $i$’s information with respect to the Blackwell order. In the remainder of this paper we will say that $F$ is partially implementable with respect to $\{\sigma_{i}\}_{i\in I}$ if truth-telling is an equilibrium in the direct mechanism in each realization of the signals $\{\sigma_{i}\}_{i\in I}$. Similarly define fully implementable with respect to $\{\sigma_{i}\}_{i\in I}.$ Note that when agents receive a signal $\sigma_{i}$ about other agents, both $\theta_{i}$ and $\sigma_{i}$ are both part of agent $i$’s private information, so if $\sigma_{i}$ has possible realizations in $\zeta_{i}=\{A_{i1},\dots,A_{ik}\}$ an implementing mechanism should ask each agent $i$ to specify a strategy for each $(\theta_{i},A_{ij})$ pair. Our setup is implicitly analyzing an equivalent formulation of this mechanism in which the uncertainty coming from the signals $\{\sigma_{i}\}$ is resolved first and asking whether $F$ is partially or fully implementable in each “interim game” of the full mechanism.
Our setup makes several substantive assumptions. Other than requiring all $\sigma_{i}$ to have full support, we assume there are no restrictions over the structure of the signal agent $i$ can receive. Thus, agents do not have private information that can constrain the types of signals they can process. Additionally, we assume the signals are not costly in order to focus on the role of the additional information regarding other agents’ types. Finally, the structure of the signals $\{\sigma_{i}\}_{i\in I}$ is assumed to be common knowledge.
2.1 The full-support signal assumption
The restriction that each $\sigma_{i}$ has full support is a simplifying assumption that avoid complications to the type space $\Theta$ after agents receive their signals. To illustrate we make this assumption, consider the following example under which agents receive perfectly informative signals. Suppose the environment consists of three agents $i=1,2,3$, where each agent $i$ has types $\theta_{i},\theta_{i}^{\prime}.$ There is a uniform common prior, so all states are equally likely. There are 5 possible outcomes $A=\{a,b,c,d,e\}.$ The SCF $F$ is as follows.
$F\mid\theta_{3}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $a$
$\theta_{1}^{\prime}$ $a$ $b$
$F\mid\theta_{3}^{\prime}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $c$ $c$
$\theta_{1}^{\prime}$ $c$ $b$
This notation specifies $F$ by separating the types of agent 1 along rows, agent 2’s types along columns, and agent 3’s types along the left versus right boxes. For example, $F(\theta_{1},\theta_{2},\theta_{3})=a$ while $F(\theta_{1},\theta_{2},\theta_{3}^{\prime})=c.$ The agents have the following preferences.
Type Preferences
$\theta_{1}$ $a\succ b\succ c\succ d\succ e$
$\theta_{1}^{\prime}$ $d\succ c\succ b\succ a\succ e$
$\theta_{2}$ $c\succ a\succ b\succ d\succ e$
$\theta_{2}^{\prime}$ $b\succ a\succ c\succ d\succ e$
$\theta_{3}$ $e\succ a\succ b\succ c\succ d$
$\theta_{3}^{\prime}$ $e\succ c\succ b\succ a\succ d$
Further, suppose agent 1’s vNM utility function $u_{1}$ is such that:
$u_{1}(a,\theta_{1})=3,u_{1}(b,\theta_{1})=2,u_{1}(c,\theta_{1})=0,$
and
$u_{1}(c,\theta_{1}^{\prime})=3,u_{1}(b,\theta_{1}^{\prime})=1,u_{1}(a,\theta_{1}^{\prime})=0.$
Then $F$ does not satisfy incentive compatibility in this information structure, since
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=\frac{3}{2}<\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1})\mid\theta_{1})=\frac{7}{4},$
and
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=\frac{5}{4}<\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=\frac{3}{2}.$
If $\sigma_{i}$ is uninformative for all $i$, it follows that $F$ is not fully or partially implementable under the specified information structure. On the other hand, if $\sigma_{i}$ is perfectly informative for all $i$, each agent $i$ has complete information regarding the state $\theta$ upon observing the realization of $\sigma_{i}$ and learning their own type $\theta_{i}$. Then incentive compatibility becomes trivial when $n\geq 3$. Each agent’s private information now consists of the complete state $\theta$. Thus if one agent unilaterally deviates from reporting their private information truthfully, the planner can simply ignore the deviation and implement the outcome in the state reported by $n-1$ agents. In this case, full implementability of $F$ reduces to the classical Nash implementation problem solved by Maskin (1999), which proves that Maskin monotonicity [^5] An SCF $F$ is Maskin-monotonic if and only if for all $\theta,\theta^{\prime}\in\Theta$, whenever $F(\theta)\neq F(\theta^{\prime})$, there exist $i\in I$ and $a\in A$ such that $u_{i}(F(\theta),\theta)\geq u_{i}(a,\theta)$ but $u_{i}(F(\theta),\theta^{\prime})<u_{i}(a,\theta^{\prime})$. and no veto power [^6] An SCF $F$ satisfies no veto power if for all $(a,j,\theta)\in A\times I\times\Theta$, the condition $u_{i}(a,\theta)\geq u_{i}(b,\theta)$ for all $b\in A$ and $i\neq j$ implies $a=F(\theta).$ are sufficient conditions for the full implementability of $F$ when there are at least three agents. No veto power is automatically satisfied because $F$ never allows two agents to simultaneously get their most preferred outcome. To see that $F$ is Maskin-monotonic, first note that outcomes $a$ and $c$ fall relative to each other in agent 3’s preference ranking when moving between types $\theta_{3}$ and $\theta_{3}^{\prime}$ and vice versa. Additionally, $b$ falls relative to $a$ in agent 1’s preference ranking when moving from type $\theta_{1}^{\prime}$ to $\theta_{1}$, and $a$ and $c$ both fall relative to $d$ when agent 1 moves from type $\theta_{1}$ to $\theta_{1}^{\prime}.$ Finally, $a$ and $c$ fall relative to each other in agent 2’s preference ranking when moving between types $\theta_{2}$ and $\theta_{2}^{\prime}$ and vice versa, and $b$ falls relative to $a$ and $c$ when moving from $\theta_{2}^{\prime}$ to $\theta_{2}.$ Thus $F$ is implementable if each $\sigma_{i}$ provides agent $i$ with a perfect signal regarding $\Theta_{-i}.$
This example demonstrates that agents receiving information about the preferences of others can allow an initially non-implementable SCF to become implementable. It also shows how the incentive compatibility condition changes when agents are able to completely rule out certain states of the world. In particular, since implementation with complete information is a special case of Bayesian implementation, it must be that $F$ satisfies incentive compatibility under complete information even though it is not incentive-compatible under the uniform common prior. This is because an agent’s type $\theta_{i}$ encapsulates information about both the preferences of agent $i$ and the information she has about the types of other agents. In the example above, $\theta_{i}$ is a complete description of agent $i$’s type under the uniform common prior, because it specifies her preferences over the outcomes but she is not able to rule out any type profile of the other agents. Under complete information, however, $\theta_{i}$ is no longer an adequate description of her type. Since agent $i$ is able to completely discern the types of the other agents, her type includes what she knows to be the type profile of the other agents. As such, if one were to translate the example above with complete information into the Bayesian implementation setup, then $\Theta_{i}=\prod_{j=1}^{3}\{\theta_{j},\theta_{j}^{\prime}\}$ for all $i.$ So, when agent $i$ has preferences according to $\theta_{i},$ she has four corresponding types $t_{i}\in\Theta_{i}$. This is consistent with the mechanism exhibited in Maskin (1999), where each agent was asked to report the complete state of the world, which included information about the preferences of all agents.
In the translation of a complete information implementation problem into one of Bayesian implementation, the states $\Theta=\prod_{i}\Theta_{i}$ contain many zero-probability states. In fact, any $t=(t_{1},t_{2},t_{3})\in\Theta$ for which each $t_{1}=t_{2}=t_{3}$ does not hold is a state that occurs with zero probability. However, the planner still specifies the SCF for these events in order to preclude equilibria that are deceptions, which trivializes incentive-compatibility when $n\geq 3$. As such, the restriction that each $\sigma_{i}$ always has full support simply prevents changes to any type space $\Theta_{i}$ after agents receive the signal realization.
3 Partial implementation and the direct mechanism
In this section we show that providing additional information to agents about the types of other agents is not helpful for the planner interested only in the partial implementation of an SCF. In our framework, this means that if $F$ is not implementable with respect to an uninformative signal structure, then it is not implementable with respect to any signal structure $\{\sigma_{i}\}_{i\in I}$. For the remainder of this paper, write $\phi_{i}$ to mean an uninformative signal about $\Theta_{-i}$ for each $i.$
Given any SCF $F$ we can consider the direct mechanism defined by $\Gamma_{F}=\langle M,h\rangle,$ where $M_{i}=\Theta_{i}$ for each $i$ and $h=F.$ Each agent $i$ has a truth-telling strategy $s:\Theta_{i}\to\Theta_{i}$ given by $s(\theta_{i})=\theta_{i}.$ The revelation principle says that $F$ is partially implementable with respect to $\{\phi_{i}\}$ iff for each agent $i\in I$ and $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$,
$\sum_{\theta_{-i}\in\Theta_{-i}}\mu(\theta_{-i}\mid\theta_{i})u_{i}(F(\theta_{i},\theta_{-i}),\theta_{i})\geq\sum_{\theta_{-i}\in\Theta_{-i}}\mu(\theta_{-i}\mid\theta_{i})u_{i}(F(\theta_{i}^{\prime},\theta_{-i}),\theta_{i}).$
If, however, $\sigma_{i}\in\{A_{i1},\dots,A_{ik}\}$ contained information about $\Theta_{-i}$, partial implementation requires $F$ to be incentive-compatible with respect to the beliefs of agent $i$ in every realization of $\sigma_{i}.$ In other words, for every $A_{i\ell}$ and $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ it must be that
$\sum_{\theta_{-i}\in\Theta_{-i}}\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=A_{i\ell})u_{i}(F(\theta_{i},\theta_{-i}),\theta_{i})\geq\sum_{\theta_{-i}\in\Theta_{-i}}\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=A_{i\ell})u_{i}(F(\theta_{i}^{\prime},\theta_{-i}),\theta_{i}).$
The law of total expectation implies that  (3) is just a weighted average of the equations  (3) for $1\leq\ell\leq k.$ This implies the following proposition.
Proposition 1.
Suppose $F$ is an SCF that is partially implementable with respect to some signal structure $\{\sigma_{i}\}_{i\in I}$. Then $F$ is partially implementable with respect to $\{\phi_{i}\}_{i\in I}$.
In other words, it can only become harder for $F$ to satisfy incentive compatibility if an agent receives an informative signal about other agents’ types, compared to the no information case. We wish to avoid ruling out the equilibrium specified by $F$ while making certain agents’ information structures more informative. Thus for the remainder of this paper we focus on the information of agents for which truth-telling is a dominant strategy in the direct mechanism.
It is well-known that the direct mechanism generally has undesired equilibria. The mechanism exhibited by Jackson (1991) to fully implement an SCF $F$ allows each agent to report, among other information, an alternate SCF that can be used to signal a deception to the planner, provided $F$ is Bayesian monotonic. In this paper, it will be useful to keep track of which agents have an incentive to alert the planner of a given deception, so we introduce the following notion.
Definition 4.
Let $F$ be an SCF and $\alpha$ be a deception such that $F\circ\alpha\neq F$. Say that $\alpha$ is undermined by agent $i$ (or agent $i$ undermines $\alpha$) if there is a function $y:\Theta_{-i}\to A$ such that
$\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}),\theta_{i})\mid\theta_{i})\geq\mathbb{E}(u_{i}(y(\theta_{-i}),\theta_{i})\mid\theta_{i})$
for all $\theta_{i}\in\Theta_{i}$ and
$\mathbb{E}(u_{i}(F(\alpha(\theta_{-i},\theta_{i}^{\prime})),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})<\mathbb{E}(u_{i}(y(\alpha_{-i}(\theta_{-i})),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})$
for some $\theta_{i}^{\prime}\in\Theta_{i}.$
The inequalities are exactly the ones from the definition of Bayesian monotonicity. Since a deception is equivalent to a strategy profile in the direct mechanism, the only deceptions $\alpha$ such that $F\circ\alpha\neq F$ that need to be undermined in order to guarantee the full implementation of $F$ are (untruthful) equilibria of the direct mechanism. For any incentive-compatible SCF $F$, write $\mathcal{A}_{F}$ to be the set of deceptions $\alpha$ such that $F\circ\alpha\neq F$ and constitute an equilibrium of the direct mechanism when agents do not have additional information about other agents. Let $\mathcal{A}_{F}^{u}\subset\mathcal{A}_{F}$ be the set of deceptions that can be undermined by some agent $i,$ again when agents have no additional information about other agents. Then we have the following.
Lemma 1.
Let $F$ be an incentive-compatible SCF. In economic environments, $F$ is fully implementable if and only if $\mathcal{A}_{F}^{u}=\mathcal{A}_{F}$.
The lemma is a direct implication of the full Bayesian implementation result of Jackson (1991) but will prove useful in our general results.
4 Main example
In this section we exhibit an SCF that is not fully implementable with respect to $\{\phi_{i}\}_{i\in I}$ but is fully implementable with respect to some informative signal structure $\{\sigma_{i}\}_{i\in I}.$ The key constraint considered in this section is Bayesian monotonicity. In particular, we show how a signal about the types of other agents can incentivize an agent to alert the planner of a deception, which the agent would not have an incentive to do without this information.
The environment has 3 agents, who each have two types, drawn non-independently from a full-support common prior over the eight possible agent type profiles. Under the planner’s desired SCF, which we will first show is incentive-compatible, there are two agents who each have an incentive to misreport her type given that the other is as well. Therefore, there is an equilibrium of the direct mechanism in which these two agents are simultaneously deceiving. Next we show that this equilibrium arises from a deception that violates Bayesian monotonicity, implying that any mechanism that partially implements the SCF will have this undesirable equilibrium. Thus the SCF is not fully implementable when agents have no additional information about other agents’ types. However, there is a third agent for which truth-telling is a dominant strategy. We show that increasing the information she has about the types of the other two agents incentivizes her to alert the planner about this deception, implying that the SCF can satisfy Bayesian monotonicity under a more informative information structure. Finally, we show that this modification to the third agent’s information allows the SCF to become fully implementable.
We now present the example formally, which will illustrate the main points of the general result. Suppose the environment consists of three agents $i=1,2,3$, where each agent $i$ has types $\theta_{i},\theta_{i}^{\prime}.$ There are 6 possible outcomes $A=\{a,b,c,d,e,f\}.$ The SCF $F$ is as follows.
$F\mid\theta_{3}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $b$
$\theta_{1}^{\prime}$ $e$ $d$
$F\mid\theta_{3}^{\prime}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $c$
$\theta_{1}^{\prime}$ $e$ $d$
The agents have the following preferences over $A$.
Type Preferences
$\theta_{1}$ $e\,(5.5)\succ a\,(4)\succ b,c\,(3)\succ d\,(2)\succ f\,(1)$
$\theta_{1}^{\prime}$ $b,c\,(5.5)\succ d\,(4)\succ e\,(3)\succ a\,(2)\succ f\,(1)$
$\theta_{2}$ $f\,(15)\succ b,c\,(4.5)\succ a\,(3)\succ e\,(2)\succ d\,(1)$
$\theta_{2}^{\prime}$ $f\,(15)\succ e\,(4.5)\succ d\,(3)\succ b,c\,(2)\succ a\,(1)$
$\theta_{3}$ $a\,(5)\succ b\,(4)\succ e\,(3)\succ d\,(2)\succ c,f\,(1)$
$\theta_{3}^{\prime}$ $a\,(5)\succ c\,(4)\succ e\,(3)\succ d\,(2)\succ b,f\,(1)$
Each agent is indifferent between two outcomes and has strict preferences over the others. The numbers in parentheses represent the utilities of each type of each agent from the outcomes.
Let $\mu$ be the common prior on $\Theta$ that places $1/6$ probability on each of the type profiles $(\theta_{1},\theta_{2},\cdot)$, $(\theta_{1}^{\prime},\theta_{2}^{\prime},\cdot),$ and places $1/3$ probability on each of the type profiles $(\theta_{1}^{\prime},\theta_{2},\cdot)$ and $(\theta_{1},\theta_{2}^{\prime},\cdot).$ Independently, each of the type profiles $(\cdot,\cdot,\theta_{3}),(\cdot,\cdot,\theta_{3}^{\prime})$ has a $1/2$ probability. In other words, the type of agent 3 is independent of the types of agents 1 and 2, but type $\theta_{1}$ of agent 1 infers $1:2$ odds of agent 2 being type $\theta_{2}$ vs. type $\theta_{2}^{\prime}$, while type $\theta_{1}^{\prime}$ of agent 1 infers $2:1$ odds of agent 2 being type $\theta_{2}$ vs. type $\theta_{2}^{\prime}$.
The outcome $f$ exists to make this environment economic, so incentive compatibility and Bayesian monotonicity are necessary and sufficient for full implementation.
The claim is that $F$ is partially but not fully implementable with respect to $\{\tau_{i}\}$ if $\tau_{i}$ is the uninformative signal for all $i.$ In particular $F$ is incentive-compatible, but there is a deception that creates a violation of Bayesian monotonicity, meaning that it cannot be ruled out by any mechanism.
4.1 Incentive compatibility
To see that $F$ is incentive-compatible, truth-telling must be an equilibrium of the direct mechanism. For agent 1, observe
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=\frac{1}{3}u_{1}(a,\theta_{1})+\frac{1}{3}u_{1}(b,\theta_{1})+\frac{1}{3}u_{1}(c,\theta_{1})=10/3$
$\geq\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1})\mid\theta_{1})=\frac{1}{3}u_{1}(e,\theta_{1})+\frac{2}{3}u_{1}(d,\theta_{1})=19/6.$
Similarly,
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=10/3\geq\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=19/6.$
Therefore, the incentive compatibility constraint holds for agent 1. A similar argument shows the incentive compatibility constraint holds for agent 2. For agent 3, truth-telling is a dominant strategy, as type $\theta_{3}$ strictly prefers outcome $b$ to $c$ while type $\theta_{3}^{\prime}$ strictly prefers outcome $c$ to $b.$ Thus, $F$ is incentive-compatible.
4.2 Bayesian monotonicity
To show that $F$ is partially but not fully implementable, we need to exhibit a deception $\alpha$ such that $F\circ\alpha\neq F$ that violates Bayesian monotonicty. We claim that $\alpha=((\theta_{1}^{\prime},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}),(\theta_{3},\theta_{3}^{\prime}))$ creates a violation of Bayesian monotonicity. In order for this to be true, it must be an equilibrium of the direct mechanism but fail the Bayesian monotonicity constraint described above. To see that $\alpha$ is an undesired equilibrium of the direct mechanism, suppose agent 1 deviates to $(\theta_{1}^{\prime},\theta_{1}).$ Then type $\theta_{2}$ of agent 2 gets an expected payoff of
$\frac{2}{3}u_{2}(a,\theta_{2})+\frac{1}{3}u_{2}(e,\theta_{2})=8/3$
from truth-telling and an expected payoff of
$\frac{1}{3}u_{2}(b,\theta_{2})+\frac{1}{3}u_{2}(c,\theta_{2})+\frac{1}{3}u_{2}(d,\theta_{2})=10/3$
from deceiving to $\theta_{2}^{\prime}$, so the best response of type $\theta_{2}$ is to deceive. Similarly type $\theta_{2}^{\prime}$ will deceive to $\theta_{2}.$ Thus $(\theta_{2}^{\prime},\theta_{2})$ is agent 2’s best response to $(\theta_{1}^{\prime},\theta_{1}).$ Similarly, $(\theta_{1}^{\prime},\theta_{1})$ is agent 1’s best response to $(\theta_{2}^{\prime},\theta_{2})$. It follows that $\alpha=((\theta_{1}^{\prime},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}),(\theta_{3},\theta_{3}^{\prime}))$ is an undesired equilibrium in the direct mechanism.
It remains to show that $\alpha$ violates Bayesian monotonicity, implying that agents 1 and 2 will always have an incentive to simultaneously deceive in any incentive-compatible mechanism. Recall that $F$ satisfies Bayesian monotonicity iff for all deceptions $\alpha$ such that $F\circ\alpha\neq F$ there exist $i\in I$ and a function $y:\Theta_{-i}\to A$ such that
$\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}),\theta_{i})\mid\theta_{i})\geq\mathbb{E}(u_{i}(y(\theta_{-i}),\theta_{i})\mid\theta_{i})$
for all $\theta_{i}\in\Theta_{i}$ and
$\mathbb{E}(u_{i}(F(\alpha(\theta_{-i},\theta_{i}^{\prime})),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})<\mathbb{E}(u_{i}(y(\alpha_{-i}(\theta_{-i})),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})$
for some $\theta_{i}^{\prime}\in\Theta_{i}.$
It needs to be shown that for no agent $i$ does there exist an alternate SCF $y$ with the desired properties. First consider agent 1. In order to define a $y:\Theta_{-1}\to A$ such that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1})\mid\theta_{1})>\mathbb{E}(u_{1}(F(\alpha(\theta_{-1},\theta_{1})),\theta_{1})\mid\theta_{1})$
it must be that (1) $y(\theta_{2},\cdot)=e$, $y(\theta_{2}^{\prime},\cdot)\succeq_{\theta_{1}}d$, and at least one of $y(\theta_{2}^{\prime},\theta_{3})\succeq_{\theta_{1}}b,c$ or $y(\theta_{2}^{\prime},\theta_{3}^{\prime})\succeq_{\theta_{1}}b,c$ holds, or (2) the following inequalities hold.
$u_{1}(y(\theta_{2},\theta_{3}),\theta_{1})+u_{1}(y(\theta_{2},\theta_{3}^{\prime}),\theta_{1})<11$
and
$u_{1}(y(\theta_{2}^{\prime},\theta_{3}),\theta_{1})+u_{1}(y(\theta_{2}^{\prime},\theta_{3}^{\prime}),\theta_{1})>4.$
If (1) applies, compute that
$\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1})\geq\frac{21}{6}>\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=\frac{10}{3},$
which is not a valid alternate SCF $y.$ Now, suppose (2) applies. Let
$A=u_{1}(y(\theta_{2},\theta_{3}),\theta_{1})+u_{1}(y(\theta_{2},\theta_{3}^{\prime}),\theta_{1}),$
and
$B=u_{1}(y(\theta_{2}^{\prime},\theta_{3}),\theta_{1})+u_{1}(y(\theta_{2}^{\prime},\theta_{3}^{\prime}),\theta_{1}).$
Then case (2) says that $A<11$ and $B>4.$ Given the utility specifications for agent 1, it follows that $A\leq 9.5$ and $B\geq 5.$ Observe that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1})\mid\theta_{1})>\mathbb{E}(u_{1}(F(\alpha(\theta_{-1},\theta_{1})),\theta_{1})\mid\theta_{1})\Rightarrow 2A+B>26,$
and
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})\geq\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1})\Rightarrow A+2B\leq 20.$
From here it is easy to see that there are no solutions to both inequalities satisfying $A\leq 9.5$ and $B\geq 5$. By similar reasoning, there is no $y:\Theta_{-1}\to A$ such that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})>\mathbb{E}(u_{1}(F(\alpha(\theta_{-1},\theta_{1}^{\prime})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})$
and
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})<\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime}).$
Thus, agent 1 cannot undermine $\alpha$. For agent 2, the utility experienced from outcome $f$ is too high for her to include it as a possible outcome in any alternate SCF $y.$ Thus, similar reasoning as what was provided for agent 1 shows that agent 2 also cannot undermine $\alpha$. For agent 3, note that her expected utility for truth-telling versus the deception $\alpha$ will always be the same, no matter what the SCF is. Therefore, the expected utility for agent 3 in the specified SCF is the same for truth-telling versus the deception $\alpha$, and similarly for any alternate SCF $y$. It follows that agent 3 can’t profit from signaling the deception via an alternate SCF unless she profits from the alternate SCF without deception as well, so agnet 3 cannot undermine $\alpha$.  [^7] This property may appear to render this example a knife-edge case, but we show in Appendix A that this example can be modified in a straightforward way to remove its reliance on agent 3’s indifference. For now, agent 3’s indifference allows the example to be somewhat simpler and focus on the effect of an additional signal. Therefore $F$ is not fully implementable if no agents receive informative signals about the types of other agents.
4.3 Eliminating $\alpha$ as an equilibrium
We claim that there is a binary signal that agent 3 can receive regarding the types of agents 1 and 2 that allows $F$ to satisfy Bayesian monotonicity and as result become fully implementable. Suppose $\sigma_{3}\in\{0,1\}$ is such that
$P(\sigma_{3}=0\mid\theta_{-3}=(\theta_{1},\theta_{2}))=P(\sigma_{3}=0\mid\theta_{-3}=(\theta_{1}^{\prime},\theta_{2}))=2/3$
and
$P(\sigma_{3}=1\mid\theta_{-3}=(\theta_{1},\theta_{2}^{\prime}))=P(\sigma_{3}=1\mid\theta_{-3}=(\theta_{1}^{\prime},\theta_{2}^{\prime}))=2/3.$
If $\sigma_{3}=0$, then agent 3’s updated beliefs place $2/9$ probability on both $(\theta_{1},\theta_{2})$ and $(\theta_{1},\theta_{2}^{\prime})$, $1/9$ probability on $(\theta_{1}^{\prime},\theta_{2}^{\prime})$, and $4/9$ probability on $(\theta_{1}^{\prime},\theta_{2}).$ If $\sigma_{3}=1,$ then agent 3’s updated beliefs place $1/9$ probability on $(\theta_{1},\theta_{2}),$ $2/9$ probability on both $(\theta_{1}^{\prime},\theta_{2})$ and $(\theta_{1}^{\prime},\theta_{2}^{\prime})$, and $4/9$ probability on $(\theta_{1},\theta_{2}^{\prime}).$
Since showing that none of the $\alpha_{i}$ for $1\leq i\leq 5$ violate Bayesian monotonicity did not require any information about agent 3’s beliefs, and the signal agent 3 receives modifies only her beliefs, it suffices to consider only the deviation $\alpha$ when showing that $F$ satisfies Bayesian monotonicity. Note that $F$ needs to satisfy Bayesian monotonicity in either realization of $\sigma_{3}.$
First suppose $\sigma_{3}=0$. Define $y:\Theta_{-3}\to A$ by $y(\theta_{1},\theta_{2})=d,y(\theta_{1},\theta_{2}^{\prime})=b,y(\theta_{1}^{\prime},\theta_{2})=e,$ and $y(\theta_{1}^{\prime},\theta_{2}^{\prime})=a.$ Then
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})=\frac{2}{9}u_{3}(a,\theta_{3})+\frac{2}{9}u_{3}(b,\theta_{3})+\frac{4}{9}u_{3}(e,\theta_{3})+\frac{1}{9}u_{3}(d,\theta_{3})=32/9$
$\geq\mathbb{E}(u_{3}(y(\theta_{-3}),\theta_{3})\mid\theta_{3})=\frac{1}{9}u_{3}(a,\theta_{3})+\frac{2}{9}u_{3}(b,\theta_{3})+\frac{4}{9}u_{3}(e,\theta_{3})+\frac{2}{9}u_{3}(d,\theta_{3})=29/9$
and
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}^{\prime}),\theta_{3}^{\prime})\mid\theta_{3}^{\prime})=32/9\geq\mathbb{E}(u_{3}(y(\theta_{-3}),\theta_{3}^{\prime})\mid\theta_{3}^{\prime})=23/9$
but
$\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3})),\theta_{3})\mid\theta_{3})=\frac{1}{9}u_{3}(a,\theta_{3})+\frac{4}{9}u_{3}(b,\theta_{3})+\frac{2}{9}u_{3}(e,\theta_{3})+\frac{2}{9}u_{3}(d,\theta_{3})=31/9$
$<\mathbb{E}(u_{3}(y(\alpha(\theta_{-3})),\theta_{3})\mid\theta_{3})=\frac{2}{9}u_{3}(a,\theta_{3})+\frac{4}{9}u_{3}(b,\theta_{3})+\frac{2}{9}u_{3}(e,\theta_{3})+\frac{1}{9}u_{3}(d,\theta_{3})=34/9.$
Now suppose $\sigma_{3}=1.$ Define $y:\Theta_{-3}\to A$ by $y(\theta_{1},\theta_{2})=a,y(\theta_{1},\theta_{2}^{\prime})=e,y(\theta_{1}^{\prime},\theta_{2})=b,y(\theta_{1}^{\prime},\theta_{2}^{\prime})=d.$ Then
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})=\frac{1}{9}u_{3}(a,\theta_{3})+\frac{4}{9}u_{3}(b,\theta_{3})+\frac{2}{9}u_{3}(e,\theta_{3})+\frac{2}{9}u_{3}(d,\theta_{3})=31/9$
$\geq\mathbb{E}(u_{3}(y(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})=\frac{1}{9}u_{3}(a,\theta_{3})+\frac{2}{9}u_{3}(b,\theta_{3})+\frac{4}{9}u_{3}(e,\theta_{3})+\frac{2}{9}u_{3}(d,\theta_{3})=29/9$
and
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}^{\prime}),\theta_{3}^{\prime})\mid\theta_{3}^{\prime})=31/9\geq\mathbb{E}(u_{3}(y(\theta_{-3},\theta_{3}^{\prime}),\theta_{3}^{\prime})\mid\theta_{3}^{\prime})=23/9.$
But
$\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3}),\theta_{3}))\mid\theta_{3})=\frac{2}{9}u_{3}(a,\theta_{3})+\frac{2}{9}u_{3}(b,\theta_{3})+\frac{4}{9}u_{3}(e,\theta_{3})+\frac{1}{9}u_{3}(d,\theta_{3})=32/9$
$<\mathbb{E}(u_{3}(y(\alpha(\theta_{-3},\theta_{3}),\theta_{3}))\mid\theta_{3})=\frac{2}{9}u_{3}(a,\theta_{3})+\frac{4}{9}u_{3}(b,\theta_{3})+\frac{2}{9}u_{3}(e,\theta_{3})+\frac{1}{9}u_{3}(d,\theta_{3})=34/9.$
Thus, upon receiving the signal $\sigma_{3}$, agent 3 can undermine $\alpha$. The intuition of the signal construction is as follows. For $F$ as described, truth-telling is a dominant strategy for agent 3. Additionally, agent 3 always prefers that agent 1 and 2’s type profile is $(\theta_{1},\theta_{2})$ rather than $(\theta_{1}^{\prime},\theta_{2}^{\prime})$, regardless of her type. Suppose agent 3 is quite confident that agent 2 is type $\theta_{2},$ so she believes it is more likely that the type profile of agents 1 and 2 belongs to $\{(\theta_{1},\theta_{2}),(\theta_{1}^{\prime},\theta_{2})\}$ rather than $\{(\theta_{1},\theta_{2}^{\prime}),(\theta_{1}^{\prime},\theta_{2}^{\prime})\}$. In particular, suppose she is confident enough that she places a higher probability on $(\theta_{1},\theta_{2})$ than $(\theta_{1}^{\prime},\theta_{2}^{\prime}).$ Then, under the SCF $F$, if the deception $\alpha$ is being played, she believes it is more likely that her utility will be harmed by agents $(\theta_{1},\theta_{2})$ deviating to $(\theta_{1}^{\prime},\theta_{2}^{\prime}),$ than benefited by the vice versa happening. It is therefore beneficial for her to propose the alternate SCF $y:\Theta_{-3}\to A$, which is defined by $y(\theta_{1},\theta_{2})=d,y(\theta_{1},\theta_{2}^{\prime})=b,y(\theta_{1}^{\prime},\theta_{2})=e,$ and $y(\theta_{1}^{\prime},\theta_{2}^{\prime})=a$, that swaps the outcomes for the type profiles $(\theta_{1},\theta_{2})$ and $(\theta_{1}^{\prime},\theta_{2}^{\prime})$ while holding the outcomes for the other type profiles constant. Similarly, agent 3 prefers that agent 1 and 2’s type profile is $(\theta_{1},\theta_{2}^{\prime})$ rather than $(\theta_{1}^{\prime},\theta_{2}).$ If agent 3 is quite confident that the type profile of agents 1 and 2 belongs to $\{(\theta_{1},\theta_{2}^{\prime}),(\theta_{1}^{\prime},\theta_{2}^{\prime})\}$, then if the deception $\alpha$ is being played with the SCF $F$, she will believe it is more likely that her utility will be harmed by agents $(\theta_{1},\theta_{2}^{\prime})$ deviating to $(\theta_{1}^{\prime},\theta_{2})$ than benefited by the vice versa happening. Thus, swapping the outcomes in these cases by proposing the alternate SCF $y:\Theta_{-3}\to A$ defined by $y(\theta_{1},\theta_{2})=a,y(\theta_{1},\theta_{2}^{\prime})=e,y(\theta_{1}^{\prime},\theta_{2})=b,$ and $y(\theta_{1}^{\prime},\theta_{2}^{\prime})=d$ allows her undermine $\alpha.$
4.4 Full implementability of $F$
We want to conclude that $F$ is implementable with respect to $\{\sigma_{i}\}_{i\in\{1,2,3\}}$ where $\sigma_{1},\sigma_{2}$ are uninformative and $\sigma_{3}$ is as specified above. To do so, we will analyze the equilibria of the direct mechanism for $F$ under this modified information structure for agent 3 and apply lemma 1.
It was shown above that truth-telling and $\alpha$ were equilibria of the direct mechanism under the uninformative information structure. It is straightforward to verify that these equilibria are unchanged in either realization of $\sigma_{3}$. The other equilibria of the direct mechanism, in either realization of $\sigma_{3}$, are the following: $\alpha_{1}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}^{\prime})),\alpha_{2}=((\theta_{1},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}^{\prime}),(\theta_{3},\theta_{3}^{\prime})),\alpha_{3}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}^{\prime})),\alpha_{4}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}))$, and $\alpha_{5}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}))$.
To see this, suppose agent 1 deviates to $(\theta_{1}^{\prime},\theta_{1}^{\prime}).$ Since both types of agent 2 prefer $e$ to $d$, agent 2’s best response is $(\theta_{2},\theta_{2}).$ Both types of agent 1 prefer $e$ to $a$, so $(\theta_{1}^{\prime},\theta_{1}^{\prime})$ is also a best response to $(\theta_{2},\theta_{2}).$ Thus $\alpha_{1}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}^{\prime}))$ is an undesired equilibrium in the direct mechanism. Similarly, if agent 1 deviates to $(\theta_{1},\theta_{1}),$ both types of agent 2 prefer $b$ or $c$ to $a,$ so agent 2’s best response is $(\theta_{2}^{\prime},\theta_{2}^{\prime}).$ Since both types of agent 1 prefer $b$ or $c$ to $d,$ agent 1’s best response is $(\theta_{1},\theta_{1}),$ so $\alpha_{2}=((\theta_{1},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}^{\prime}),(\theta_{3},\theta_{3}^{\prime}))$ is another undesired equilibrium in the direct mechanism. Finally, since agent 3’s report only affects the outcome of the direct mechanism if agent 1 reports $\theta_{1}$ and agent 2 reports $\theta_{2}^{\prime}$, it follows that $\alpha_{3}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}^{\prime})),\alpha_{4}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}))$, and $\alpha_{5}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}))$ are all undesired equilibria, in either realization of $\sigma_{3}$. Truth-telling, $\alpha,\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},$ and $\alpha_{5}$ cover all possible strategies for agent 1 in the direct mechanism. Since they are all equilibria and truth-telling is a dominant strategy for agent 3 (a property that is independent of $\sigma_{3}$’s realization), there can be no other equilibria in the direct mechanism, in either realization of $\sigma_{3}$.
We claim that each deception $\alpha_{i}$ for $1\leq i\leq 5$ can be undermined by either agent 1 or agent 2. To see this, first note that all such deceptions make the expected utility of agents 1 and 2 certain, since agents 1 and 2 are not sending different messages for different realizations of their own type. Moreover, in no case does a deception $\alpha_{i}$ allow all types of agents 1 and 2 to for sure get their most preferred outcome. This makes it easy for the agents to define an alternate SCF that would benefit them. For example, in the case of $\alpha_{1},$ consider the function $y:\Theta_{-1}\to A$ defined by $y(\theta_{2},\cdot)=d,y(\theta_{2}^{\prime},\cdot)=f$. Then
$E(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=10/3\geq E(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1})=4/3$
and
$E(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=10/3\geq E(u_{1}(y(\theta_{-1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=3$
but
$E(u_{1}(y(\alpha_{1}(\theta_{-1})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=4>E(u_{1}(F(\alpha_{1}(\theta_{-1},\theta_{1}^{\prime})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=3.$
It follows that $\alpha_{1}$ is undermined by agent 1. Similar arguments show that $\alpha_{2},\alpha_{3},\alpha_{4},$ and $\alpha_{5}$ are also undermined. Since $\sigma_{3}$ only affects the information of agent 3, each $\alpha_{i}$ for $1\leq i\leq 5$ can be undermined by either agent 1 or agent 2 in both realizations of $\sigma_{3}$. In the previous subsection we saw that agent 3 can undermine $\alpha$ in either realization. Therefore, lemma 1 implies that $F$ is fully implementable with respect to $\{\sigma_{i}\}_{i\in\{1,2,3\}}$.
This example also illustrates how a planner can strictly prefer agents to have an information structure that is between no information and complete information. Indeed, $F$ is also not fully implementable in a complete information environment because it fails Maskin monotonicity: when moving from the state $(\theta_{1}^{\prime},\theta_{2},\theta_{3})$ to $(\theta_{1},\theta_{2},\theta_{3})$, the outcome $e$ is chosen in the former state and weakly rises in every agent’s preference ordering, but $e$ is not chosen in the latter state.
5 General results
The main example suggests how an agent for which truth-telling is a dominant strategy for a given SCF $F$ can undermine a given deception $\alpha$ with additional information about the types of other agents, in a way that she would not be able to do so without this information. In particular, suppose $F$ is an SCF such that there is a type profile $\theta_{-i}^{\prime}$ that yields an outcome agent $i$ prefers to that of the type profiles in $\alpha_{-i}(\theta_{-i}^{\prime})$, regardless of what agent $i$ reports. Suppose additionally that agent $i$ believes $\theta_{-i}^{\prime}$ is more likely to be the true type profile of other agents than the type profiles in $\alpha_{-i}^{-1}(\theta_{-i}^{\prime})$. Then it is possible for some type $\theta_{i}$ of agent $i$ to undermine $\alpha$ by proposing the alternate SCF that swaps the outcomes corresponding to $\theta_{-i}^{\prime}$ and $\alpha_{-i}(\theta_{-i}^{\prime})$ assuming she plays $\alpha_{i}(\theta_{i})$. Thus, if agent $i$ can receive a signal that ensures such a type profile $\theta_{-i}^{\prime}$ always exists, then an undesirable equilibrium $\alpha$ can be potentially undermined. We formalized this into the following condition.
Definition 5.
Let $F$ be an SCF and $\alpha\in\mathcal{A}_{F}$. Agent $i$ can potentially undermine $\alpha$ if truth-telling is a dominant strategy for agent $i$ and there exist distinct type profiles $\theta_{-i}^{\prime},\theta_{-i}^{\prime\prime}$ such that $\alpha_{-i}(\alpha_{-i}(\theta_{-i}^{\prime}))\neq\theta_{-i}^{\prime\prime}$, $\alpha_{-i}(\alpha_{-i}(\theta_{-i}^{\prime\prime}))\neq\theta_{-i}^{\prime}$, $\alpha_{-i}(\theta_{-i}^{\prime})\neq\alpha_{-i}(\theta_{-i}^{\prime\prime})$ and the inequalities
$u_{i}(F(\theta_{-i}^{\prime},\theta_{i}^{\prime}),\theta_{i}^{\prime})\geq u_{i}(F(\alpha_{-i}(\theta_{-i}^{\prime}),\theta_{i}^{\prime}),\theta_{i}^{\prime}),$
$u_{i}(F(\theta_{-i}^{\prime\prime},\theta_{i}^{\prime}),\theta_{i}^{\prime})\geq u_{i}(F(\alpha_{-i}(\theta_{-i}^{\prime\prime}),\theta_{i}^{\prime}),\theta_{i}^{\prime})$
hold for all $\theta_{i}^{\prime}\in\Theta_{i}.$ Furthermore the inequalities are strict for some type $\theta_{i}^{\prime}=\theta_{i}.$
The inequalities capture the intuition that agent $i$ should always prefer other agents report truthfully rather than according to the deception $\alpha$ if she knew the type profiles $\theta_{-i}^{\prime}$ or $\theta_{-i}^{\prime\prime}$ realized. This preference needs to be strict for at least one type of agent $i$ in order to provide sufficient incentives for her to deviate from the undesirable equilibrium $\alpha.$ The requirements that $\alpha_{-i}(\alpha_{-i}(\theta_{-i}^{\prime}))\neq\theta_{-i}^{\prime\prime},\alpha_{-i}(\alpha_{-i}(\theta_{-i}^{\prime\prime}))\neq\theta_{-i}^{\prime},$ and $\alpha_{-i}(\theta_{-i}^{\prime})\neq\alpha_{-i}(\theta_{-i}^{\prime\prime}),$ limit the interactions $\theta_{-i}^{\prime}$ and $\theta_{-i}^{\prime\prime}$ have with each other in the deception $\alpha$, which allows us to partition the type space $\Theta_{-i}$ in a natural way. As the name suggests, an agent $i$ that can potentially undermine a deception $\alpha$ can undermine $\alpha$ with additional information about the types of other agents. In other words, we have the following lemma, which is the main technical contribution of this paper.
Lemma 2.
Let $F$ be an incentive-compatible SCF $\alpha$ be a deception that agent $i$ can potentially undermine. Then there is a binary signal $\sigma_{i}:\Theta_{-i}\to\Delta(\{0,1\})$ such that agent $i$ can undermine $\alpha$ after receiving either realization of $\sigma_{i}.$
In particular, for $\theta_{-i}^{\prime},\theta_{-i}^{\prime\prime}$ as in Definition 5, the signal informs agent $i$, with sufficiently high accuracy, whether or not the type profile of other agents is in the set
$\{\theta_{-i}^{\prime},\alpha_{-i}(\theta_{-i}^{\prime\prime})\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{\prime}).$
The full proof is given in Appendix B. For an incentive-compatible $F$ that is not fully implementable, the lemma suggests a systematic way to eliminate all deceptions in $\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}$ (which must be non-empty by lemma 1) with additional information. In particular, we have the following result.
Theorem 1.
Let $F$ be an incentive-compatible SCF. If there is a subset $J\subset I$ such that each $\alpha\in\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}$ can be potentially undermined by a distinct agent $j\in J$, and every $\alpha\in\mathcal{A}_{F}^{u}$ can be undermined by some agent $i\in I\setminus J$, then $F$ is fully implementable with respect to some signal structure $\{\sigma_{i}\}_{i\in I}.$
The result is essentially a straightforward application of lemmas 1 and 2, and the proof is in the appendix. The theorem requires a separation of the agents who can undermine deceptions in $\mathcal{A}_{F}^{u}$ from the agents who can potentially undermine deceptions in $\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}$ because we cannot guarantee that an agent who undermines a deception $\alpha\in\mathcal{A}_{F}^{u}$ without additional information can still undermine $\alpha$ after receiving an informative signal about other agents. This implicitly requires $|\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}|\leq n-1$.
6 Conclusion
The main finding of the paper is that increasing the information agents have about the types of other agents can serve to eliminate undesired equilibria in full implementation problems. Consequently, there are social choice rules that are not originally fully implementable under a given information structure but become fully implementable after agents receive noisy signals about the types of other agents. We provide a sufficient condition that guarantees the existence of a more informative information structure that can allow an SCF $F$ that is not fully implementable under some full-support common prior to become fully implementable. We conclude by discussing some limitations of the analysis.
The results are dependent on the specifics of agents’ beliefs in different information structures. In particular, given a non-fully-implementable SCF, any increase in information agents have about the types of other agents would need to be common knowledge for there to potentially exist a mechanism that fully implements the SCF under the new information structure. While these assumptions are standard in the implementation literature, a recent strand of literature in mechanism design focuses on the robustness of mechanisms (pioneered by Bergemann and Morris (2005)), arguing that mechanism design should not rely excessively on common knowledge assumptions among the agents and the planner. The dependence on common knowledge assumptions is certainly a limitation on the results presented, but it is worth noting the main finding — that increasing information an agent has about other agents can eliminate undesired equilibria — does not depend on a single, specific change in information. Rather, in every example and theorem presented, the signal an agent receives that eliminates an undesired equilibrium is simply required to be above a certain level of accuracy. In all of the examples and results presented, whenever a signal $\sigma$ of accuracy $10.5<\tau<$ 1 eliminates an undesired equilibrium, any signal with the same realizations but higher accuracy $\tau^{\prime}>\tau$ also eliminates the equilibrium. Thus, while the agents do need to fully understand when other agents have received additional information, the exact accuracy of the signals received does not need to be known.
Appendix A Appendix A: A more robust example
In the example provided in section 4, agent 3 is indifferent between the truth-telling equilibrium and the deception $\alpha$ regardless of the social choice rule, raising the question of whether it is simply describing a knife-edge case. But the example can be modified to accommodate this concern.
There are still three agents $i=1,2,3$, where each agent $i$ has types $\theta_{i},\theta_{i}^{\prime},$ and 6 possible outcomes $A=\{a,b,c,d,e,f\}.$ The SCF $F$ remains the same, as follows.
$F\mid\theta_{3}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $b$
$\theta_{1}^{\prime}$ $e$ $d$
$F\mid\theta_{3}^{\prime}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $c$
$\theta_{1}^{\prime}$ $e$ $d$
Modify the agents’ preferences to get the following.
Type Preferences
$\theta_{1}$ $e\,(6)\succ a\,(4)\succ b,c\,(3)\succ d\,(2)\succ f\,(1)$
$\theta_{1}^{\prime}$ $b,c\,(6)\succ d\,(4)\succ e\,(3)\succ a\,(2)\succ f\,(1)$
$\theta_{2}$ $f\,(25)\succ b,c\,(7)\succ a\,(3)\succ e\,(2)\succ d\,(1)$
$\theta_{2}^{\prime}$ $f\,(20)\succ e\,(4.5)\succ d\,(3)\succ b,c\,(2)\succ a\,(1)$
$\theta_{3}$ $a\,(5)\succ b\,(4)\succ f\,(3)\succ c\,(2)\succ d,e\,(1)$
$\theta_{3}^{\prime}$ $a\,(5)\succ c\,(4)\succ f\,(3)\succ b\,(2)\succ d,e\,(1)$
Modify the common prior $\mu$ to be defined by
$\mu(\theta_{1},\theta_{2},\cdot)=1/10,\mu(\theta_{1},\theta_{2}^{\prime},\cdot)=3/10,\mu(\theta_{1}^{\prime},\theta_{2},\cdot)=9/20,\mu(\theta_{1}^{\prime},\theta_{2}^{\prime},\cdot)=3/20.$
Moreover, the type of agent 3 is independent of the types of agents 1 and 2, and $\mu(\cdot,\cdot,\theta_{3})=\mu(\cdot,\cdot,\theta_{3}^{\prime})=1/2$.
We follow similar steps as in the main example to show that $F$ is incentive-compatible and that the direct mechanism contains the equilibria $\alpha=((\theta_{1}^{\prime},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}),(\theta_{3},\theta_{3}^{\prime}))$, $\alpha_{1}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}^{\prime})),\alpha_{2}=((\theta_{1},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}^{\prime}),(\theta_{3},\theta_{3}^{\prime})),\alpha_{3}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}^{\prime})),\alpha_{4}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}))$, and $\alpha_{5}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}))$. It is also analogous to show that all of the $\alpha_{i}$ for $1\leq i\leq 5$ can be undermined by agents 1 and 2.
The deception $\alpha$ still violates Bayesian monotonicity under the specified common prior, meaning that no agent can undermine $\alpha$. As before, first consider agent 1. In order to define a $y:\Theta_{-1}\to A$ such that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1})\mid\theta_{1})>\mathbb{E}(u_{1}(F(\alpha_{-1}(\theta_{-1},\theta_{1})),\theta_{1})\mid\theta_{1})$
it must be that (1) $y(\theta_{2},\cdot)=e$, $y(\theta_{2}^{\prime},\cdot)\succeq_{\theta_{1}}d$, and at least one of $y(\theta_{2}^{\prime},\theta_{3})\succeq_{\theta_{1}}b,c$ or $y(\theta_{2}^{\prime},\theta_{3}^{\prime})\succeq_{\theta_{1}}b,c$ holds; (2) $y(\theta_{2},\theta_{3})=e$, $y(\theta_{2},\theta_{3}^{\prime})=a$ and $y(\theta_{2}^{\prime},\cdot)\succeq_{\theta_{1}}b,c$; or (3) $y(\theta_{2},\theta_{3})=a$, $y(\theta_{2},\theta_{3}^{\prime})=e$ and $y(\theta_{2}^{\prime},\cdot)\succeq_{\theta_{1}}b,c$. If (1) applies, compute that
$\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1})\geq\frac{27}{8}>\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=\frac{13}{4},$
which is not a valid alternate SCF $y.$ If any of the other cases apply, it must also be that
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})<\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1}).$
By similar reasoning, in order to define a $y:\Theta_{-1}\to A$ such that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})>\mathbb{E}(u_{1}(F(\alpha(\theta_{-1},\theta_{1}^{\prime})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})$
it must also be that
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})<\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime}).$
Thus agent 1 cannot undermine $\alpha$. For agent 2, the utility experienced from outcome $f$ is again too high for her to include it as a possible outcome in any alternate SCF $y,$ so agent 2 cannot undermine $\alpha$. Since agent 3 is no longer indifferent between $\alpha$ and truth-telling, it remains to show that no $y:\Theta_{-3}\to A$ exists that makes $F$ Bayesian monotonic. Note that
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})=\frac{1}{10}u_{3}(a,\theta_{3})+\frac{3}{10}u_{3}(b,\theta_{3})+\frac{9}{20}u_{3}(e,\theta_{3})+\frac{3}{20}u_{3}(d,\theta_{3})$
while
$\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3})),\theta_{3})\mid\theta_{3})=\frac{3}{20}u_{3}(a,\theta_{3})+\frac{9}{20}u_{3}(b,\theta_{3})+\frac{3}{10}u_{3}(e,\theta_{3})+\frac{1}{10}u_{3}(d,\theta_{3}).$
Then there are 2 ways to define a $y:\Theta_{-3}\to A$ to get
$\mathbb{E}(u_{3}(y(\alpha(\theta_{-3})),\theta_{3})\mid\theta_{3})>\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3})),\theta_{3})\mid\theta_{3}).$
An alternate SCF $y$ must either satisfy (1) $y(\theta_{1},\cdot)=a$ and $y(\theta_{1}^{\prime},\cdot)\in\{d,e\}$ or (2) $y(\theta_{1}^{\prime},\theta_{2})\succ_{\theta_{3}}d,e$ or $y(\theta_{1}^{\prime},\theta_{2}^{\prime})\succ_{\theta_{3}}d,e$. If (1) applies, it is clear that
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})<\mathbb{E}(u_{3}(y(\theta_{-3}),\theta_{3})\mid\theta_{3}).$
Now suppose (2) applies. Then in order for
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})\geq\mathbb{E}(u_{3}(y(\theta_{-3}),\theta_{3})\mid\theta_{3})$
it must be that
$u_{3}(a,\theta_{3})+3u_{3}(b,\theta_{3})-u_{3}(y(\theta_{1},\theta_{2}),\theta_{3})-3u_{3}(y(\theta_{1},\theta_{2}^{\prime}),\theta_{3})$
$\geq\frac{3}{2}\left(3u_{3}(y(\theta_{1}^{\prime},\theta_{2}),\theta_{3})+u_{3}(y(\theta_{1}^{\prime},\theta_{2}^{\prime}),\theta_{3})-3u_{3}(e,\theta_{3})-u_{3}(d,\theta_{3})\right).$
On the other hand, for
$\mathbb{E}(u_{3}(y(\alpha(\theta_{-3})),\theta_{3})\mid\theta_{3})>\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3})),\theta_{3})\mid\theta_{3})$
to hold, it must be that
$3u_{3}(y(\theta_{1}^{\prime},\theta_{2}),\theta_{3})+u_{3}(y(\theta_{1}^{\prime},\theta_{2}^{\prime}),\theta_{3})-3u_{3}(e,\theta_{3})-u_{3}(d,\theta_{3})$
$>\frac{3}{2}(u_{3}(a,\theta_{3})+3u_{3}(b,\theta_{3})-u_{3}(y(\theta_{1},\theta_{2}),\theta_{3})-3u_{3}(y(\theta_{1},\theta_{2}^{\prime}),\theta_{3})).$
Since all quantities are positive, it is impossible for (iii) and (iv) to hold simultaneously. Thus there cannot be a $y:\Theta_{-3}\to A$ that allows $F$ to satisfy Bayesian monotonicity.
Similar reasoning as what was given in the main example implies the signal $\sigma_{3}\in\{0,1\}$ such that
$P(\sigma_{3}=0\mid\theta_{-3}=(\theta_{1},\theta_{2}))=P(\sigma_{3}=0\mid\theta_{-3}=(\theta_{1}^{\prime},\theta_{2}))=2/3$
and
$P(\sigma_{3}=1\mid\theta_{-3}=(\theta_{1},\theta_{2}^{\prime}))=P(\sigma_{3}=1\mid\theta_{-3}=(\theta_{1}^{\prime},\theta_{2}^{\prime}))=2/3,$
if accessed by agent 3, allows $F$ to satisfy Bayesian monotonicity.
Appendix B Appendix B: Omitted proofs
Proof of Proposition 1.
Proof.
Fix a agent $j.$ Since every agents’ beliefs and every signal has full support, the only possible signal that affects agent $j$’s expected payoff is $\sigma_{j}.$ Suppose $\sigma_{j}\in\{A_{j1},\dots,A_{jk}\}$ and let $q_{\ell}$ be the unconditional probability that $\sigma_{j}=A_{j\ell}$ (so $q_{\ell}$ is averaged over all possible states). If $F$ is incentive-compatible with respect to agent $j$ after agent $j$ receives this signal, then
$q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{j\ell})\geq q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{j\ell})$
for all $1\leq\ell\leq k.$ By the law of iterated expectations,
$\sum_{\ell=1}^{k}q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{j\ell})=\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j}),$
and
$\sum_{\ell=1}^{k}q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{j\ell})=\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j}).$
Therefore, it must be that
$\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j})\geq\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j}),$
so $F$ was incentive-compatible prior to agent $j$ receiving $\sigma_{j}$. ∎
Proof of Lemma 1.
Proof.
The only if direction is direct from the definitions. For the if direction, consider the following mechanism $\Gamma=\langle M,h\rangle$ to implement the SCF $F.$ The mechanism, slightly adapted from Jackson (1991) to implement a single SCF $F$ rather than a set of SCFs, was shown to fully implement any SCF $F$ satisfying incentive compatibility and Bayesian monotonicity in economic environments. Let $S=\max_{i}|\Theta_{i}|$ and $K=n(S+1).$ Let $V=\{0,1,\dots,S^{2}\}^{K},$ so $V$ consists of $K$-dimensional integer vectors with entries between 00 and $S^{2}.$ Let $X$ be the space of all possible SCFs $F:\Theta\to A.$ Let
$M_{i}=\Theta_{i}\times\{\{\varnothing\}\cup V\}\times X\times\{\{\varnothing\}\cup X\}.$
For any message profile $m=(m_{1},\dots,m_{n})\in M$, let $\theta^{m}\in\Theta$ be the type profile given by the first entry of every agent’s message. The following rules specify $h$.
Rule 1: Suppose there is a SCR $F\in\widehat{F}$ and $i\in N$ such that every $m_{j}$ for $j\neq i$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ). Additionally, $m_{i}$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ), or (,,,).($\cdot$,$\cdot$,$\cdot$,$\varnothing$).( ⋅ , ⋅ , ⋅ , ∅ ) . Then $h(m)=F(\theta^{m}).$
Rule 2: Suppose there is a SCR $F\in\widehat{F}$ and $i\in N$ such that every $m_{j}$ for $j\neq i$ is of the form (,,,)($\cdot$,$\varnothing$,$\cdot$,$\varnothing$)( ⋅ , ∅ , ⋅ , ∅ ). Additionally, $m_{i}$ is of the form $(\cdot,\cdot,\cdot,y)$ for some $y\in X$. Let $\theta_{i}$ be the first entry of $m_{i},$ and define $y_{\theta_{i}}\in X$ by $y_{\theta_{i}}(\theta)=y(\theta_{-i},\theta_{i})$ for all $\theta\in\Theta.$ Then
$h(m)=\begin{cases}y(\theta^{m})&\text{if }\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})\geq\mathbb{E}(u_{i}(y_{\theta_{i}}(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})\text{ for all }\theta_{i}^{\prime}\in\Theta_{i}\\ F(\theta^{m})&\text{if }\mathbb{E}(u_{i}(y_{\theta_{i}}(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})>\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime})\text{ for some }\theta_{i}^{\prime}\in\Theta_{i}\end{cases}.$
Rule 3: Suppose neither rule 1 nor rule 2 applies. For $1\leq k\leq 4,$ let $m_{i}^{k}$ denote the $k$th entry of the message of agent $i.$ Then $h(m)=m_{i^{*}}^{3}(\theta^{m})$, where $i^{*}$ is determined by the following. Let $I_{0}$ be the set of agents $i\in I$ that choose to specify an integer vector in their message, so $I_{0}=\{i\mid m_{i}^{2}\neq\varnothing\}$. For $i\in I_{0}$ denote $m_{i}^{2}$ by $v^{i},$ and let the $l$th coordinate of the vector $v^{i}$ be $v_{l}^{j}.$ Let
$J(i)=\#\{j\in I_{0}\mid v_{\ell}^{i}=v_{i}^{j}\text{ for an integer }\ell\text{ such that }n+(j-1)S<\ell\leq n+jS\}.$
If there is $i$ such that $J(i)>J(k)$ for all $k\in I_{0}$ and $k\neq i$, then $i^{*}=i.$ Otherwise, $i^{*}=\min_{i}(I_{0}).$
Rules 2 and 3 ensure that in an equilibrium $s$ of $\Gamma$, Rule 1 must apply, as otherwise one agent can unilaterally deviate to get her most preferred outcome. If rule 1 applies and $s$ is an equilibrium of $\Gamma,$ it must be the case that $(s_{i}^{*})_{i\in I}$ — where $s_{i}^{*}(\theta_{i})$ is the first coordinate of $s_{i}(\theta_{i})$ for all $i$ and $\theta_{i}\in\Theta_{i}$ — is an equilibrium of the direct mechanism. Rule 2 ensures that $s=(s_{i})_{i\in I}$ where $s_{i}(\theta_{i})$ is of the form $(\theta_{i},\varnothing,\cdot,\varnothing)$ for all $i$ and $\theta_{i}\in\Theta_{i}$ is a BNE of $\Gamma$, so $\Gamma$ clearly partially implements $F.$ If $\mathcal{A}_{F}=\mathcal{A}_{F}^{u}$ then rule 2 ensures that no $s$ such that $(s_{i}^{*})_{i\in I}$ is an undesirable equilibrium of the direct mechanism constitutes an equilibrium of $\Gamma.$ Thus $\Gamma$ fully implements $F$ whenever $\mathcal{A}_{F}=\mathcal{A}_{F}^{u}$. ∎
Proof of Lemma 2.
Proof.
Suppose agent $i$ can potentially undermine $\alpha$. Therefore truth-telling is a dominant strategy for agent $i$ and there exist a type $\theta_{i}\in\Theta_{i}$ as well as distinct type profiles $\theta_{-i}^{1},\theta_{-i}^{2},\theta_{-i}^{3},\theta_{-i}^{4}\in\Theta_{-i}$ such that $\alpha_{-i}(\theta_{-i}^{4})\neq\theta_{-i}^{1},\alpha_{-i}(\theta_{-i}^{2})\neq\theta_{-i}^{3}$, and $\alpha_{-i}(\theta_{-i}^{j})=\theta_{-i}^{j+1}$ for $j=1,3$. Furthermore,
$u_{i}(F(\theta_{-i}^{j},\theta_{i}^{\prime}),\theta_{i}^{\prime})\geq u_{i}(F(\theta_{-i}^{j+1},\theta_{i}^{\prime}),\theta_{i}^{\prime})$
for $j=1,3$ and all $\theta_{i}^{\prime}\in\Theta_{i},$ and the inequalities are strict for type $\theta_{i}^{\prime}=\theta_{i}.$ Let $\sigma_{i}\in\{0,1\}$ be a binary signal such that if $\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3})$, then $P(\sigma_{i}=0\mid\theta_{-i})=\tau$ for some accuracy $10.5<\tau<$ 1. Additionally, if $\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3})$, then $P(\sigma_{i}=1\mid\theta_{-i})=\tau.$ With slight abuse of notation we write $\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}$ to mean $\{\theta_{-i}^{1},\theta_{-i}^{4}\}\cup\alpha_{-i}^{-1}(\theta_{-i}^{3}).$ Since agent $i$ can potentially undermine $\alpha$, we have from the definition that $\{\alpha_{-i}^{-1}(\theta_{-i}^{1}),\theta_{-i}^{2},\theta_{-i}^{3}\}\subset\Theta_{-i}\setminus\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}$. Thus, upon receiving $\sigma_{i}=0$, agent $i$ will place probability
$\tau\mu(\theta_{-i}^{\prime}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})$
on all $\theta_{-i}^{\prime}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}$ and probability
$(1-\tau)\mu(\theta_{-i}^{\prime}\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})$
on all $\theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}.$ Here $\mu(\cdot\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})$ represents the beliefs agent $i$ would have over $\Theta_{-i}$ conditional on knowing her type is $\theta_{i}^{\prime}$, and the type profile $\theta_{-i}$ of other agents is in the set $\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}.$ We will show that there exists an accurate enough $\tau$ such that agent $i$ can undermine $\alpha$.
To prove the lemma, $\alpha$ must be undermined by agent $i$ in either realization of the signal, so first suppose $\sigma_{i}=0.$ Now let $y:\Theta_{-i}\to A$ be defined by $y(\theta_{-i}^{1})=F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),y(\theta_{-i}^{2})=F(\theta_{-i}^{1},\alpha_{i}(\theta_{i}))$, and $y(\theta_{-i})=F(\theta_{-i},\alpha_{i}(\theta_{i}))$ for all $\theta_{-i}\neq\theta_{-i}^{1},\theta_{-i}^{2}.$ This construction of $y$ is a general formulation of the alternate SCFs constructed for the main example. It first needs to be shown that
$\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\sigma_{i}=0)\geq\mathbb{E}(u_{i}(y(\theta_{-i}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\sigma_{i}=0)$
for all $\theta_{i}^{\prime}\in\Theta_{i}.$ To see this, observe that the following holds for sufficiently accurate $\tau$.
$\mathbb{E}(u_{i}(F$ $(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\sigma_{i}=0)=\tau\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})$
$+(1-\tau)\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})$
$\geq\tau\mu(\theta_{-i}^{1}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\theta_{i}^{\prime}),\theta_{i}^{\prime})$
$+\tau\mu(\theta_{-i}^{4}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\theta_{i}^{\prime}),\theta_{i}^{\prime})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{\prime},\theta_{i}^{\prime}),\theta_{i}^{\prime})$
$+(1-\tau)\mu(\theta_{-i}^{2}\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{1},\theta_{i}^{\prime}),\theta_{i}^{\prime})$
$+(1-\tau)\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\in\Theta_{-i}\setminus\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\\ \theta_{-i}^{\prime}\neq\theta_{-i}^{2}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{\prime},\theta_{i}^{\prime}),\theta_{i}^{\prime})$
$\geq\tau\mu(\theta_{-i}^{1}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i}^{\prime})$
$+\tau\mu(\theta_{-i}^{4}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i}^{\prime})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{\prime},\alpha_{i}(\theta_{i})),\theta_{i}^{\prime})$
$+(1-\tau)\mu(\theta_{-i}^{2}\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{1},\alpha_{i}(\theta_{i})),\theta_{i}^{\prime})$
$+(1-\tau)\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\in\Theta_{-i}\setminus\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\\ \theta_{-i}^{\prime}\neq\theta_{-i}^{2}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{\prime},\alpha_{i}(\theta_{i})),\theta_{i}^{\prime})$
$=\mathbb{E}(u_{i}(y(\theta_{-i}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\sigma_{i}=0),$
as desired. The first inequality is implied by our assumption that agent $i$ always prefers $\theta_{-i}^{1}$ to $\theta_{-i}^{2}$ and holds whenever
$\tau\mu(\theta_{-i}^{1}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})\geq(1-\tau)\mu(\theta_{-i}^{2}\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}),$
which can be done for sufficiently large $\tau.$ The second inequality is because truth-telling is a dominant strategy for agent $i.$ The last equality follows from the definition of $y.$ Thus, the desired inequality holds.
It also needs to be shown that there is sufficiently large $\tau$ such that
$\mathbb{E}(u_{i}(F(\alpha(\theta_{-i},\theta_{i})),\theta_{i})\mid\theta_{i},\sigma_{i}=0)<\mathbb{E}(u_{i}(y(\alpha_{-i}(\theta_{-i})),\theta_{i})\mid\theta_{i},\sigma_{i}=0).$
To see this, as before let $\alpha_{-i}^{-1}(\theta_{-i}^{1})$ be the (possibly empty) set of $\theta_{-i}$ such that $\alpha_{-i}(\theta_{-i})=\theta_{-i}^{1}$. Note that our assumptions imply that $\alpha_{-i}^{-1}(\theta_{-i}^{1})$ is disjoint from $\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}$. So,
$\mathbb{E}(u_{i}$ $(F(\alpha(\theta_{-i},\theta_{i})),\theta_{i})\mid\theta_{i},\sigma_{i}=0)=\tau\mu(\theta_{-i}^{1}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\mu(\theta_{-i}^{4}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{4}),\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{1})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{1},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\notin\alpha_{-i}^{-1}(\theta_{-i}^{1})\\ \theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{\prime}),\alpha_{i}(\theta_{i})),\theta_{i}).$
Additionally,
$\mathbb{E}(u_{i}$ $(y(\alpha_{-i}(\theta_{-i})),\alpha_{i}(\theta_{i}))\mid\theta_{i},\sigma_{i}=0)=\tau\mu(\theta_{-i}^{1}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{1},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\mu(\theta_{-i}^{4}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{4}),\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{1})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\notin\alpha_{-i}^{-1}(\theta_{-i}^{1})\\ \theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(y(\alpha_{-i}(\theta_{-i}^{\prime})),\theta_{i}).$
Comparing these expressions, first observe that
$\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\notin\alpha_{-i}^{-1}(\theta_{-i}^{1})\\ \theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{\prime}),\alpha_{i}(\theta_{i})),\theta_{i})$
$\leq\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\notin\alpha_{-i}^{-1}(\theta_{-i}^{1})\\ \theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(y(\alpha_{-i}(\theta_{-i}^{\prime})),\theta_{i}).$
This is because the only possible $\theta_{-i}$ for which terms in the sum will differ are for $\theta_{-i}\in\alpha_{-i}^{-1}(\theta_{-i}^{2})$ such that $\theta_{-i}\neq\theta_{-i}^{1}.$ In the former summation, the corresponding term will be
$K_{2}=\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i}).$
The corresponding term in the latter summation, which is
$K_{1}=\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{1},\alpha_{i}(\theta_{i})),\theta_{i}).$
If $\alpha_{i}(\theta_{i})=\theta_{i},$ then $K_{2}<K_{1}$ is automatic from the requirement that agent $i$ can potentially undermine $\alpha$. If $\alpha_{i}(\theta_{i})\neq\theta_{i}$, then $u_{i}$ must satisfy
$u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i})=u_{i}(F(\theta_{-i}^{2},\theta_{i}),\theta_{i})$
and
$u_{i}(F(\theta_{-i}^{1},\alpha_{i}(\theta_{i})),\theta_{i})=u_{i}(F(\theta_{-i}^{1},\theta_{i}),\theta_{i})$
because truth-telling is a dominant strategy for agent $i.$ Thus the assumption that agent $i$ can potentially undermine $\alpha$ also implies $K_{2}<K_{1}.$
Thus, to guarantee
$\mathbb{E}(u_{i}(F(\alpha(\theta_{-i},\theta_{i})),\theta_{i})\mid\theta_{i},\sigma_{i}=0)<\mathbb{E}(u_{i}(y(\alpha_{-i}(\theta_{-i})),\theta_{i})\mid\theta_{i},\sigma_{i}=0),$
it suffices to pick $\tau$ sufficiently large so that
$\tau\mu(\theta_{-i}^{1}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{1})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{1},\alpha_{i}(\theta_{i})),\theta_{i})$
$<\tau\mu(\theta_{-i}^{1}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{1},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{1})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i}),$
which is again possible by the condition that agent $i$ can potentially undermine $\alpha$ and similar reasoning as above. Thus $\alpha$ is undermined in the signal realization $\sigma_{i}=0.$
If the signal realization is $\sigma_{i}=1,$ one can similarly define $y:\Theta_{-i}\to A$ by $y(\theta_{-i}^{3})=F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),y(\theta_{-i}^{4})=F(\theta_{-i}^{3},\alpha_{i}(\theta_{i}))$, and $y(\theta_{-i})=F(\theta_{-i},\alpha_{i}(\theta_{i}))$ for all $\theta_{-i}\neq\theta_{-i}^{3},\theta_{-i}^{4}.$ The proof that $\alpha$ is undermined in this case is largely analogous. To show that
$\mathbb{E}(u_{i}(F(\theta_{-i},\theta_{i}^{\prime}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\sigma_{i}=1)\geq\mathbb{E}(u_{i}(y(\theta_{-i}),\theta_{i}^{\prime})\mid\theta_{i}^{\prime},\sigma_{i}=1)$
for all $\theta_{i}^{\prime}\in\Theta_{i}$, we can use similar steps as above to show that $\tau$ needs to satisfy
$\tau\mu(\theta_{-i}^{3}\mid\theta_{i}^{\prime},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})\geq(1-\tau)\mu(\theta_{-i}^{4}\mid\theta_{i}^{\prime},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}),$
for all $\theta_{i}^{\prime}\in\Theta_{i}.$
Finally, one again needs to check that there is sufficiently large $\tau$ such that
$\mathbb{E}(u_{i}(F(\alpha(\theta_{-i},\theta_{i})),\theta_{i})\mid\theta_{i},\sigma_{i}=1)<\mathbb{E}(u_{i}(y(\alpha_{-i}(\theta_{-i})),\theta_{i})\mid\theta_{i},\sigma_{i}=1).$
Observe that
$\mathbb{E}(u_{i}$ $(F(\alpha(\theta_{-i},\theta_{i})),\theta_{i})\mid\theta_{i},\sigma_{i}=1)=(1-\tau)\mu(\theta_{-i}^{1}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\mu(\theta_{-i}^{4}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{4}),\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{4})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\notin\alpha_{-i}^{-1}(\theta_{-i}^{4})\\ \theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{\prime}),\alpha_{i}(\theta_{i})),\theta_{i}),$
while
$\mathbb{E}(u_{i}$ $(y(\alpha(\theta_{-i})),\theta_{i})\mid\theta_{i},\sigma_{i}=1)=(1-\tau)\mu(\theta_{-i}^{1}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{2},\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\mu(\theta_{-i}^{4}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{4}),\alpha_{i}(\theta_{i})),\theta_{i})$
$+(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{4})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\begin{subarray}{c}\theta_{-i}^{\prime}\notin\alpha_{-i}^{-1}(\theta_{-i}^{4})\\ \theta_{-i}^{\prime}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\}\end{subarray}}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\alpha_{-i}(\theta_{-i}^{\prime}),\alpha_{i}(\theta_{i})),\theta_{i}).$
It therefore suffices to pick $\tau$ sufficiently large so that
$(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{4})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i})$
$<(1-\tau)\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{3})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\in\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{4},\alpha_{i}(\theta_{i})),\theta_{i})$
$+\tau\sum_{\theta_{-i}^{\prime}\in\alpha_{-i}^{-1}(\theta_{-i}^{4})}\mu(\theta_{-i}^{\prime}\mid\theta_{i},\theta_{-i}\notin\{\theta_{-i}^{1},\theta_{-i}^{4},\alpha_{-i}^{-1}(\theta_{-i}^{3})\})u_{i}(F(\theta_{-i}^{3},\alpha_{i}(\theta_{i})),\theta_{i}).$
As before, this can be done with sufficiently large $\tau$. Thus, after agent $i$ receives the signal $\sigma_{i}$, she can undermine $\alpha$, and the proof is complete. ∎
Proof of Theorem 1.
Proof.
Let $\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}=\{\alpha_{1},\dots,\alpha_{k}\}$ and $j_{\ell}$ be the agent that can potentially undermine $\alpha_{\ell}$ for all $1\leq\ell\leq k$. By lemma 2, for each $\ell$ there is a binary signal $\sigma_{j_{\ell}}$ such that agent $j_{\ell}$ can undermine $\alpha_{\ell}$ upon receiving $\sigma_{j_{\ell}}$. By assumption each $j_{\ell}$ is distinct and $\{j_{1},\dots,j_{k}\}\subset J.$ For all $i\in I\setminus\{j_{1},\dots,j_{k}\}$ let $\sigma_{i}$ be uninformative. Lemma 1 directly implies that $F$ is fully implementable with respect to these signals $\{\sigma_{i}\}_{i\in I}$ as long as the set of equilibria in the direct mechanism is unchanged after the agents $j_{1},\dots,j_{k}$ receive their signals. More precisely, if we take $\mathcal{B}_{F}$ to be the set of Bayes Nash equilibria of the direct mechanism when agents only know the common prior $\mu\in\Delta(\Theta)$, we want to show that $s\in\mathcal{B}_{F}$ if and only if $s$ is an equilibrium of the direct mechanism after the agents $j_{1},\dots,j_{k}$ receive their signals.
To see this, let $s=(s_{1},\dots,s_{n})$ be a strategy profile. Then $s$ is an equilibrium in the direct mechanism when no agent has additional information about other agents iff for all $i$ and $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ we have
$\sum_{\theta_{-i}}u(F(s_{i}(\theta_{i}),s_{-i}(\theta_{-i})),\theta_{i})\mu(\theta_{-i}\mid\theta_{i})\geq\sum_{\theta_{-i}}u(F(\theta_{i}^{\prime},s_{-i}(\theta_{-i})),\theta_{i})\mu(\theta_{-i}\mid\theta_{i}).$
On the other hand, $s$ is an equilibrium in the direct mechanism after agents receive the signal realizations $\{\sigma_{i}\}_{i\in I}$ iff for all $i\in I$, $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$, and signal realizations $x$ in the range of $\sigma_{i}$, we have
$\sum_{\theta_{-i}}u(F(s_{i}(\theta_{i}),s_{-i}(\theta_{-i})),\theta_{i})\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=x)\geq\sum_{\theta_{-i}}u(F(\theta_{i}^{\prime},s_{-i}(\theta_{-i})),\theta_{i})\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=x).$
When $i\notin\{j_{1},\dots,j_{k}\}$ the signal $\sigma_{i}$ is uninformative so the inequalities don’t change between (v) and (vi). If $i\in\{j_{1},\dots,j_{k}\}$, then truth-telling is a dominant strategy for agent $i$, so it must be that
$u(F(\theta_{i},s_{-i}(\theta_{-i})),\theta_{i})\geq u(F(\theta_{i}^{\prime},s_{-i}(\theta_{-i})),\theta_{i})$
for all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$, since $\mu$ and $\sigma_{i}$ have full support. Thus both (v) and (vi) are imply
$u(F(s_{i}(\theta_{i}),s_{-i}(\theta_{-i})),\theta_{i})=u(F(\theta_{i},s_{-i}(\theta_{-i})),\theta_{i})$
for all $i\in\{j_{1},\dots,j_{k}\}$ and $\theta_{i}\in\Theta_{i}$. This completes the proof. ∎